At what distance does a star have an apparent magnitude that is 5 magnitudes brighter (smaller) than its absolute magnitude?


1
Expert's answer
2021-01-01T14:54:30-0500

Let's apply the formula

"M=m+5-5 \lg{(r)}"

Where

"M-" absolute magnitude

"m-" apparent magnitude

where do we write

"M-m=5-5 \lg{(r)}"

by problem statement

"M-m=5"

then

"5=5-5 \lg{(r)}"

"5-5=-5 \lg{(r)}"

"0=-5 \lg{(r)}"

"0=\lg{(r)}"

from where

"r=1 pc"